# Accelerated signal-to-noise ratio interpolation across the gravitational-wave signal manifold powered by artificial neural networks

### v1

@@ Abstract
Matched-filter based gravitational-wave search pipelines identify candidate events within seconds of their arrival on Earth, offering a chance to guide electromagnetic follow-up and observe multi-messenger events. Understanding the detectors’ response to an astrophysical signal across the searched signal manifold is paramount to inferring the parameters of the progenitor and deciding which candidates warrant telescope time. In this paper, we use artificial neural networks to accelerate and signal-to-noise ratio (SNR) computation for sufficiently local patches of the signal manifold. Our machine-learning based model generates a single waveform (or equivalently, computes a single SNR timeseries) in 6 milliseconds on a CPU and 0.4 milliseconds on a GPU. When we use the GPU to generate batches of waveforms simultaneously, we find that we can produce $10^{4}$ waveforms in $\lesssim 1$ ms on a GPU. This is achieved while remaining faithful, on average, to 1 part in $10^{4}$ (1 part in $10^{5}$) for binary black hole (binary neutron star) waveforms. The model we present is designed to directly utilize intermediate detection pipeline outputs, and is a step towards ultra-low-latency parameter estimation within search pipelines.
@@ I Introduction
Gravitational-waves (GW) observed by Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) have provided an unprecedented view into our Universe. The candidates identified by the LIGO-Virgo-Kagra collaboration (LVK) Abbott et al. (2021) and external groups Nitz et al. (2019); Magee et al. (2019); Venumadhav et al. (2020); Zackay et al. (2021); Nitz et al. (2020, 2021, 2023); Olsen et al. (2022) provide an increasingly complete census of the ultracompact binaries in our universe Abbott et al. (2023). Much of this science has been made possible by the low-latency search pipelines that analyze detector data in near-real-time to rapidly identify candidate GWs across a wide range of masses Tsukada et al. (2023); Nitz et al. (2018); Aubin et al. (2021); Chu et al. (2022); Drago et al. (2020).
Binary neutron stars (BNSs) are of particular interest since they are known multi-messenger sources Abbott et al. (2017a, b). The rapid and accurate localization Singer and Price (2016), classification Farr et al. (2015); Kapadia et al. (2020); Andres et al. (2022); Villa-Ortega et al. (2022), and source property estimation Foucart et al. (2018); Ghosh et al. (2021) of low-latency GW candidates clarify which events warrant electromagnetic follow-up. These processes are presently seeded by information from the single most significant candidate observed across LVK detection pipelines. Although this results in precise localizations Singer et al. (2014), incorporating information from across the entire searched parameter space will lead to a systematically better understanding of low-latency candidates Chaudhary et al. (2024).
@@
Recently, machine-learning algorithms have emerged as a promising way to navigate GW science and other areas of data-driven astrophysics. This has largely been made possible by advances to graphical processing units (GPUs) and the ease of software that utilize them NVIDIA et al. (2020); Paszke et al. (2019); xla . In GW astrophysics alone, machine-learning algorithms have already been applied to compact binary detection Schäfer et al. (2023); Mishra et al. (2022); Krastev (2020); Schäfer et al. (2020), early warning alerts Wei and Huerta (2021); Baltus et al. (2021); Yu et al. (2021); Baltus et al. (2022), supernovae identification Astone et al. (2018); Antelis et al. (2022); López Portilla et al. (2021), parameter estimation Gabbard et al. (2022); Dax et al. (2021); Chatterjee et al. (2022); Chatterjee and Wen (2022); Dax et al. (2023); Wong et al. (2023), noise removal and characterization Vajente et al. (2020); Essick et al. (2020); Saleem et al. (2023), and waveform interpolation Schmidt et al. (2021); Chua et al. (2019); Khan and Green (2021); Tissino et al. (2023); Thomas et al. (2022).
@@
Previous waveform interpolation applications have primarily focused on aligned-spin binary black holes (BBH) Chua et al. (2019); Khan and Green (2021); Schmidt et al. (2021), though some have considered non-aligned systems as well Thomas et al. (2022). Each of these works first converts the waveforms to a new basis, which is achieved via methods such as principal component analysis and Gram-Schmidt orthonormalization, before interpolating the coefficients needed to reconstruct the waveforms. In this work, we describe a similar technique, with extra attention paid to the case of interpolating intermediate detection pipeline outputs. We apply neural networks, a form of supervised machine learning, to the interpolation of the GW emission and signal-to-noise ratio (SNR) computation from arbitrary aligned-spin compact binaries using singular value decomposition (SVD). This work is inspired by previous studies that used grid-based techniques to interpolate the SVD of nonspinning BBH signals Cannon et al. (2012a), as well as mesh-free approaches applied to the SVD of aligned-spin binary neutron star (BNS) emission Pathak et al. (2022). We aim to leverage information collected across GW searches by providing a general interpolation technique that utilizes the intermediate SNR timeseries collected by detection pipelines. We build on previous SVD interpolation efforts and additionally consider the impacts of component spin, interpolate waveforms of arbitrary length, and provide a method that is immediately applicable to existing low-latency analyses and requires minimal fine tuning.
This paper is organized as follows. First, we describe the algorithms used to generate our training data. Second, we introduce the architecture of our neural network and present results for binary black hole and binary neutron star systems. Finally, we discuss future applications of this work in the direction of near-real-time inference via an analog of the reduced-order-quadrature method Canizares et al. (2015); Smith et al. (2016); Morisaki and Raymond (2020).
@@ II.1 Background and motivation
Modern matched-filter based GW pipelines simultaneously search for GWs from BNS, BBH, and neutron star black hole (NSBH) binaries. The searches precompute the expected gravitational wave emission, also known as a template, for $\mathcal{O}(10^{6})$ binaries representative of the space. This collection of templates, or template bank, is designed to capture $97\%-99\%$ of the signals that fall within the search space. While the overall parameter space is vast, pipelines often divide their search across a number of partially-overlapping sub-banks in the signal manifold that contain templates with morphologies Messick et al. (2017); Roulet et al. (2019); Sakon et al. (2022) that respond similarly to astrophysical sources and detector noise.
The similarities between the waveforms in a sub-bank can be exploited by using singular value decomposition (SVD) Cannon et al. (2010); Roulet et al. (2019) to reduce the filtering cost of the search. The SVD ensures that we can express any of the $N$ waveforms within our sub-bank as a sum of some orthogonal basis vectors, $u_{\mu}(t)$, and complex-valued reconstruction coefficients, $a(\lambda)_{\mu}$:
@@
where $\lambda=(m_{1},m_{2},s_{1z},s_{2z})$ represents the set of intrinsic parameters of the system, $h(\lambda,t)$ is the whitened gravitational-wave strain for parameters $\lambda$ at time $t$, $M$ represents the number of bases identified by the SVD, and $M\leq N$. Further approximations can be applied to ensure that $M<<N$, which greatly reduces the number of filters a search must correlate with the data. We refer the reader to Cannon et al. (2010) for a complete description of this method applied to compact binary searches.
@@ II.2 Machine learning architecture
The number of layers, neurons per layer, and choice of activation function vary wildly depending on the task to the neural network is applied to. In this work, we use the network to perform regression and estimate a set of weights that describes the contribution for a number of basis vectors. We use a neural network consisting of a 4-node input layer corresponding to $(m_{1},m_{2},s_{1z},s_{2z})$, 4 hidden layers of 512 neurons, and an output layer whose size depends on the total number of basis vectors contained in a specific sub-bank (Fig. 1). For our data, each time slice is independent of the next, so we have the freedom to fit each segment of the waveform separately. Similarly, we could train each basis vector independently. Instead, we elect to fit all time slices and bases at once to reduce the total number of models needed. We use Rectified Linear Units (ReLUs) as the activation function for our hidden layers, defined as
@@
We use the mean-square error as our loss function, and the Adam optimizer Kingma and Ba (2014) to minimize the loss function. We choose to use $\lambda$ as the input to our network to highlight the generality of this method, but we note that transforming coordinates to the parameters most important for the phasing of the signal Cutler and Flanagan (1994); Morisaki and Raymond (2020) could lead to more accurate or faster converging results.
@@ III.1 Fidelity
We apply the above formalism to BNS and BBH systems. We generate small BNS and BBH banks with minimum matches Owen and Sathyaprakash (1999) of $0.99$ using a stochastic placement algorithm Harry et al. (2009); Privitera et al. (2014). The banks are constructed in a constrained section of the parameter space to be representative of sub-bank grouping in current low-latency analyses Sakon et al. (2022). The BBH (BNS) bank is designed to recover compact binaries with $20M_{\odot}<m_{2}\leq m_{1}<40M_{\odot}$ ($1.40M_{\odot}<m_{2}\leq m_{1}<1.42M_{\odot}$) and $0.750.5<s_{iz}<$ 0.75 ($0.010.00<s_{iz}<$ 0.01), where $m_{1},m_{2}$ are the component masses of the binary and $s_{1z},s_{2z}$ are the components of the spin angular momentum aligned with the orbital angular momentum. The resulting bank contains 193 (343) waveforms. We verify the coverage of the banks by drawing $10^{4}$ random samples from within the bounds and computing the mismatches,
@@
Figure 3: Two representative interpolated waveforms and the associated residual phase. The vertical dashed line in the left plot indicates the merger time for the BBH. The pre-merger BBH phase agrees to approximately 1 part in $10^{1}-10^{2}$, while the binary neutron star phase agrees to 1 part in $10^{2}-10^{4}$. This shows that the interpolant provided here accurately reconstructs both the amplitude and phase of arbitrary signals.
We train a neural network for each bank separately using the architecture described in Section II.2. We find that both the BBH and BNS models are computed in $\lesssim 10\,\mathrm{minutes}$ on a GPU. The performance of the resulting models are summarized in Fig. 2. We find that our BBH (BNS) model achieves a median mismatch of $10^{-4}$ ($10^{-5}$). The worst mismatch is $10^{-2}$ ($10^{-3}$), which is of the order of waveform systematics Bohé et al. (2017). The worst reconstructed waveforms are shown in Figure 3a and 3b alongside their residuals for amplitude and phase.
III.2 Accelerated waveform generation and SNR calculation
@@
We find that the neural network based interpolant described here has similar performance to the above methods when utilizing accelerated linear algebra xla and just-in-time compilation. We summarize our findings in Table 1. We find that we can produce a single waveform in 0.5 ms on a CPU and 0.3 ms in a GPU. GPUs also offer the ability to batch waveform generation. On an A100 GPU, we find that we can produce $10^{4}$ waveforms in $\sim 0.8$ ms, which corresponds to an effective speed of 1 waveform per 80 ns. The waveform generation time is roughly independent of signal length since the size of the output layer size remains approximately constant across banks.
@@
Recall that under our chosen framework, generating a GW via a weighted average of SVD bases is equivalent to calculating the SNR timeseries using a weighted average of SNR timeseries collected for each orthogonal basis found by the SVD (see Equation 4). The timings described in Table 1 are thus also representative of how long it takes to calculate an arbitrary SNR using the orthogonal SNRs output by the LLOID algorithm in low-latency. For reference, we compare to the cost of performing this correlation on a CPU with Bilby Ashton et al. (2019). Table 2 shows the average time to compute $\langle h(t)|d(t)\rangle$ in the frequency domain for 16 s (64 s) BBH (BNS) waveforms at 5 different sample rates. Naively comparing our timing results to those of Bilby indicate a $1-30$ time speed-up for a single SNR calculation and an $\mathcal{O}(10^{4}-10^{5})$ speed-up for calculations batched on a GPU. This method is similar to reduced-order-quadrature (ROQ) techniques Canizares et al. (2015); Smith et al. (2016); Morisaki and Raymond (2020), which offer similar accelerations for much broader signal manifolds than the ones considered here, though these are not directly integrated into the detection process.
@@ III.3 Signal loss from PSD drift
Since the basis vectors are constructed via SVD on the whitened gravitational waveforms, changes in the measured PSD fundamentally alter the bases and can lead to signal loss. To minimize bias, the low-latency GstLAL analysis recomputes the bases on a weekly cadence with updated PSDs. While in principle, the same should be done for our method, we find it to be a small effect in practice. To estimate the impact, we use the more optimistic, publicly available aligoO4_high.txt PSD estimate for O4444This PSD is also available at and represents LIGO detectors that are $12.5\%$ more sensitive than those described by aligoO4_low.txt. This is an extreme case of PSD drift since the new PSD represents a detector that is $12.5\%$ more sensitive. We use this as a proxy for PSD evolution over the course of the run. With this new PSD, we generate $1000$ sample whitened waveforms and compare to the waveforms predicted by our neural network. We find that models trained on the less sensitive PSD still reproduce the correct waveform to within 1 part in $10^{2}$ ($10^{3}$). These results are shown in Figure 4. This is still within the range of waveform systematics, though we do not rule out the possibility that specific, frequency dependent PSD changes could lead to more substantial differences. In the case of model breaking PSD changes, our network can be asynchronously updated and retrained with the new spectrum. We leave a study of potential biases on parameter estimation to a future work.

### v2

# A neural network-based gravitational wave interpolant with applications to low-latency analyses

@@ Abstract
Matched-filter based gravitational-wave search pipelines identify candidate events within seconds of their arrival on Earth, offering a chance to guide electromagnetic follow-up and observe multi-messenger events. Understanding the detectors’ response to an astrophysical transient across the searched signal manifold is paramount to inferring the parameters of the progenitor and deciding which candidates warrant telescope time. We describe a framework that uses artificial neural networks to interpolate gravitational waves and, equivalently, the signal-to-noise ratio (SNR) across sufficiently local patches of the signal manifold. Our machine-learning based model generates a single waveform in 6 milliseconds on a CPU and 0.4 milliseconds on a GPU. When using a GPU to generate batches of waveforms simultaneously, we find that we can produce $10^{4}$ waveforms in $\lesssim 1$ ms. This is achieved while remaining faithful, on average, to 1 part in $10^{4}$ (1 part in $10^{5}$) for binary black hole (binary neutron star) waveforms. The model we present is designed to directly utilize intermediate detection pipeline outputs in the hopes of facilitating a better real-time understanding of gravitational-wave candidates.
@@ I Introduction
Gravitational-waves (GW) observed by Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) have provided an unprecedented view into our Universe. The candidates identified by the LIGO-Virgo-Kagra collaboration (LVK) Abbott et al. (2021) and external groups Nitz et al. (2019); Magee et al. (2019); Venumadhav et al. (2020); Zackay et al. (2021); Nitz et al. (2020, 2021, 2023); Olsen et al. (2022) provide an increasingly complete census of the ultracompact binaries in our universe Abbott et al. (2023). Much of this science has been made possible by the low-latency search pipelines Tsukada et al. (2023); Nitz et al. (2018); Aubin et al. (2021); Chu et al. (2022); Drago et al. (2020) that rapidly identifies GW candidates and the alert infrastructure Magee et al. (2021); Chaudhary et al. (2024) that subsequently localizes and classifies them.
Binary neutron stars (BNSs) and neutron star – black hole binaries (NSBHs) remain particularly sought after since they are known Abbott et al. (2017a, b) (or, in the case of NSBHs, expected Foucart et al. (2018)) sources of electromagnetic radiation. It is therefore crucial to precisely understand the nature of compact binaries in near-real-time to determine which warrant electromagnetic followup. This is increasingly important as highly anticipated, next generation facilities like the Rubin Observatory Ivezić et al. (2019) begin observations in the coming years. At present, our low-latency estimates of localization Singer and Price (2016), classification Farr et al. (2015); Kapadia et al. (2020); Andres et al. (2022); Villa-Ortega et al. (2022), and source properties  Foucart et al. (2018); Ghosh et al. (2021) are derived from the highest signal-to-noise ratio (SNR) event among the maximum likelihood candidates identified by LVK detection pipelines. Though this results in precise localizations Singer et al. (2014), it is well understood that full parameter estimation is systematically more accurate Chaudhary et al. (2024). Unfortunately, even in their low-latency configuration, parameter estimation pipelines can take $\mathcal{O}(\mathrm{hours})$ to produce stable results, which may be too late to facilitate the capture of rapidly fading transients Cowperthwaite et al. (2017).
@@
In this work, we examine neural networks as a way to interpolate the search response, in the hopes of developing a better real-time understanding of the nature of the candidate and enabling rapid parameter estimation in the future. Recently, machine-learning algorithms have emerged as a promising way to navigate GW science and other areas of data-driven astrophysics. This has largely been made possible by advances to graphical processing units (GPUs) and the ease of software that utilize them NVIDIA et al. (2020); Paszke et al. (2019); Sabne (2020). In GW astrophysics alone, machine-learning algorithms have already been applied to compact binary detection George and Huerta (2018); Gabbard et al. (2018); Schäfer et al. (2023); Mishra et al. (2022); Krastev (2020); Schäfer et al. (2020, 2023); Nousi et al. (2023); Marx et al. (2024), early warning alerts Wei and Huerta (2021); Baltus et al. (2021); Yu et al. (2021); Baltus et al. (2022); Alfaidi and Messenger (2024), supernovae identification Astone et al. (2018); Antelis et al. (2022); López Portilla et al. (2021), parameter estimation Gabbard et al. (2022); Dax et al. (2021); Chatterjee et al. (2022); Chatterjee and Wen (2022); Dax et al. (2023); Wong et al. (2023), noise removal and characterization Vajente et al. (2020); Essick et al. (2020); Saleem et al. (2023), and waveform interpolation Schmidt et al. (2021); Chua et al. (2019); Khan and Green (2021); Tissino et al. (2023); Thomas et al. (2022).
@@
Previous waveform interpolation applications have primarily focused on aligned-spin binary black holes (BBH) Chua et al. (2019); Khan and Green (2021); Schmidt et al. (2021), though some have considered non-aligned systems as well Thomas et al. (2022). Each of these works first converts the waveforms to a new basis, which is achieved via methods such as principal component analysis and Gram-Schmidt orthonormalization, before interpolating the coefficients needed to reconstruct the waveforms. In this work, we describe a similar technique. We employ neural networks trained via supervised learning to the interpolation of the GW emission and SNR of arbitrary aligned-spin compact binaries using singular value decomposition (SVD). This work is inspired by previous studies that used grid-based techniques to interpolate the SVD of nonspinning BBH signals Cannon et al. (2012a), as well as mesh-free approaches applied to the SVD of aligned-spin BNS emission Pathak et al. (2022). We build on previous SVD interpolation efforts and additionally consider the impacts of component spin, interpolate waveforms of arbitrary length, and provide a method that is immediately applicable to existing low-latency analyses and requires minimal fine tuning. The interpolant we describe provides a way to rapidly calculate the SNR of aligned-spin compact binaries, which may be able to accelerate low-latency data enrichment tasks.
This paper is organized as follows. First, we describe the matched-filtering algorithms and neural network architecture in Section II. Second, we apply this network to BBH and BNS systems. We describe the accuracy of the interpolant on these systems and present a timing study in Section III. Finally, we conclude with a discussion of future applications in low-latency environments in Section IV.
@@ II.1 Background and motivation
Modern matched-filter based GW pipelines simultaneously search for GWs from BNS, BBH, and NSBH binaries. The searches precompute the expected gravitational wave emission, also known as a template, for $\mathcal{O}(10^{6})$ binaries representative of the space. This collection of templates, or template bank, is designed to retain $97\%-99\%$ of the SNR. While the overall parameter space is vast, pipelines often divide their search across a number of small, partially-overlapping sub-banks in the signal manifold that contain templates with morphologies Messick et al. (2017); Roulet et al. (2019); Sakon et al. (2022) that respond similarly to astrophysical sources and detector noise.
The similarities between the waveforms in a sub-bank can be exploited by using singular value decomposition (SVD) Cannon et al. (2010); Roulet et al. (2019) to reduce the filtering cost of the search. The SVD ensures that we can express any of the $N$ [^1] Performing an SVD on large matrices is known to be computationally expensive. Within the GstLAL framework, searches set $N\leq 1024$ to reduce cost. waveforms within our sub-bank as a sum of some orthogonal basis vectors, $u_{\mu}(t)$, and complex-valued reconstruction coefficients, $a(\lambda)_{\mu}$:
@@
where $\lambda=(m_{1},m_{2},s_{1z},s_{2z})$ represents the set of intrinsic parameters of the system, $h(\lambda,t)$ is the whitened gravitational-wave strain for parameters $\lambda$ at time $t$, $M$ represents the number of bases identified by the SVD, and $M\leq N$. We neglect the impact of the location and orientation of the binary on the GW emission, which can be applied via a complex rotation for aligned-spin systems. Further approximations can be applied to ensure that $M<<N$, which greatly reduces the number of filters a search must correlate with the data. We refer the reader to Cannon et al. (2010) for a complete description of this method applied to compact binary searches.
@@ II.2 Machine learning architecture
The appropriate choice for the number of layers, neurons per layer, and activation function can vary depending on the task, though there are a number of key features that are typically desired. Continuous differentiability, monotonicity, and fixed output ranges are generally preferred since they facilitate network backpropagation, gradient descent, and learning stability, and can ameliorate the vanishing or exploding gradient problem Goodfellow et al. (2016). In this work, we use the network to perform regression and estimate a set of weights that describes the contribution for a number of basis vectors. We use a neural network consisting of a 4-node input layer corresponding to $\lambda=(m_{1},m_{2},s_{1z},s_{2z})$, 4 hidden layers of 512 neurons, and an output layer whose size depends on the total number of basis vectors contained in a specific sub-bank (Fig. 1). For our data, each time slice is independent of the next, so we have the freedom to fit each segment of the waveform separately. Similarly, we could train each basis vector independently. Instead, we elect to fit all time slices and bases at once to reduce the total number of models needed. We use Rectified Linear Units (ReLUs) as the activation function for our hidden layers, defined as
@@
We use the mean-square error as our loss function, and the Adam optimizer Kingma and Ba (2014) to minimize the loss function. We choose to use $\lambda$ as the input to our network to highlight the generality of this method, but we note that transforming coordinates to the parameters most important for the phasing of the signal Cutler and Flanagan (1994); Morisaki and Raymond (2020) could lead to more accurate or faster converging results. We do not exclude the possibility that another set of coordinates may be necessary in regions of the parameter space that exhibit significant degeneracies.
@@ III.1 Fidelity
We apply the above formalism to BNS and BBH systems. We generate small BNS and BBH banks with minimum matches Owen and Sathyaprakash (1999) of $0.99$ using a stochastic placement algorithm Harry et al. (2009); Privitera et al. (2014). The banks are not intended to be representative of a full GW search; rather, they are constructed in a constrained section of the parameter space to be illustrative of sub-bank grouping in current low-latency analyses. We refer the reader to Figures 2 and 4 in Sachdev et al. (2019) and Sakon et al. (2022), respectively, for grouping visualizations. The BBH (BNS) bank is designed to recover compact binaries with $20M_{\odot}<m_{2}\leq m_{1}<40M_{\odot}$ ($1.40M_{\odot}<m_{2}\leq m_{1}<1.42M_{\odot}$) and $0.750.5<s_{iz}<$ 0.75 ($0.010.00<s_{iz}<$ 0.01), where $m_{1},m_{2}$ are the component masses of the binary and $s_{1z},s_{2z}$ are the components of the spin angular momentum aligned with the orbital angular momentum. The resulting bank contains 193 (343) waveforms. We verify the coverage of the banks by drawing $10^{4}$ random samples from within the bounds and computing the mismatches,
@@
Figure 3: Two representative interpolated waveforms and the associated residual phase. The vertical dashed line in the left plot indicates the merger time for the BBH. The pre-merger BBH phase agrees to approximately 1 part in $10^{1}-10^{2}$, while the BNS phase agrees to 1 part in $10^{2}-10^{4}$. This shows that the interpolant provided here accurately reconstructs both the amplitude and phase of arbitrary signals.
We train a neural network for each bank separately using the architecture described in Section II.2. We find that both the BBH and BNS models are computed in $\lesssim 10\,\mathrm{minutes}$ on a GPU. The performance of the resulting models are summarized in Fig. 2. We find that our BBH (BNS) model achieves a median mismatch of $10^{-4}$ ($10^{-5}$). The worst mismatch is $10^{-2}$ ($10^{-3}$), which is of the order of waveform systematics Bohé et al. (2017). The worst reconstructed waveforms are shown in Figure 3a and 3b alongside their phase residuals.
III.2 Waveform and SNR interpolation timing study
@@
We find that the neural network based interpolant described here has similar performance to the above methods when utilizing accelerated linear algebra Sabne (2020) and just-in-time compilation. We summarize our findings in Table 1. We find that we can produce a single waveform in 0.5 ms on a CPU and 0.3 ms in a GPU. GPUs also offer the ability to batch waveform generation. On an A100 GPU, we find that we can produce $10^{4}$ waveforms in $\sim 0.8$ ms, which corresponds to an effective speed of 1 waveform per 80 ns. The waveform generation time is roughly independent of signal length since the size of the output layer size remains approximately constant across banks. In all cases, we randomly sample from within the confines of the reference template bank.
@@
Recall that under our chosen framework, generating a GW via a weighted average of SVD bases is equivalent to calculating the SNR timeseries using a weighted average of SNR timeseries collected for each orthogonal basis found by the SVD (see Equation 4). The timings described in Table 1 are thus also representative of how long it takes to calculate an arbitrary SNR using the orthogonal SNRs output by the LLOID algorithm in low-latency. For reference, we compare to the cost of performing this correlation on a CPU with Bilby Ashton et al. (2019). Table 2 shows the average time to compute $\langle h(t)|d(t)\rangle$ in the frequency domain for a fiducial 16 s (64 s) BBH (BNS) waveforms at 5 different sample rates. Naively comparing our timing results to those of Bilby indicate a $1-30$ time speed-up for a single SNR calculation and an $\mathcal{O}(10^{4}-10^{5})$ speed-up for calculations batched on a GPU. In each case, we compare the time to produce an SNR value when either Bilby or our method is evaluated at a known point in the parameter space. This method is similar to reduced-order-quadrature (ROQ) techniques Canizares et al. (2015); Smith et al. (2016); Morisaki and Raymond (2020), which offer similar accelerations for much broader signal manifolds than the ones considered here, though these are not directly integrated into the detection process.
@@ III.3 Signal loss from PSD drift
Since the basis vectors are constructed via SVD on the whitened gravitational waveforms, changes in the measured PSD fundamentally alter the bases and can lead to signal loss. To minimize bias, the low-latency GstLAL analysis recomputes the bases on a weekly cadence with updated PSDs. While in principle, the same should be done for our method, we find it to be a small effect in practice. To estimate the impact, we use the more optimistic, publicly available aligoO4_high.txt PSD estimate for O4555This PSD is also available at and represents LIGO detectors that are $12.5\%$ more sensitive than those described by aligoO4_low.txt. This is an extreme case of PSD drift since the new PSD represents a detector that is $12.5\%$ more sensitive. We use this as a proxy for PSD evolution over the course of the run. With this new PSD, we generate $1000$ sample whitened waveforms and compare to the waveforms predicted by our neural network. We find that models trained on the less sensitive PSD still reproduce the correct waveform to within 1 part in $10^{2}$ ($10^{3}$). These results are shown in Figure 4. This is still within the range of waveform systematics, though we do not rule out the possibility that specific, frequency dependent PSD changes could lead to more substantial differences. In the case of model breaking PSD changes, our network can be asynchronously updated and retrained with the new spectrum. We leave a study of potential biases on parameter estimation to a future work.